OpenStudy (anonymous):

Use an addition or subtraction formula to find the exact value. cos (-pi/12)

OpenStudy (anonymous):

cos (-pi/2) = cos (pi/2) = 0

OpenStudy (anonymous):

Or, using addition/subtraction formula, cos ( 0 - pi/2) = cos 0 cos pi/2 + sin 0 sin pi/2 = 1(0) + 0(1) = 0

OpenStudy (anonymous):

0- pi/2 doesn't give you -pi/12. @Easyaspi314

OpenStudy (anonymous):

Sorry, I read the question wrong, I thought it said -pi/2

OpenStudy (anonymous):

- pi/12 = -15 degrees

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Use the formula for cos(30-45) = cos - 15 degrees

OpenStudy (anonymous):

pi radians = 180 degrees, so - pi/12 radians = -180/12 = -15 degrees

OpenStudy (anonymous):

In your text there is a formula for cos (x - y), the difference formula for cos

OpenStudy (anonymous):

So cos (30 - 45) according to the formula = cos 30 cos 45 + sin 30 sin 45

OpenStudy (anonymous):

You're suppoised to know that cos 30 = sqrt(3)/2; cos 45 = sqrt(2)/2; sin 30 = 1/2; sin (45) = sqrt (2)/2 so cos (30-45) = what? plug in